OpenStudy (anonymous):

If the vertices of an ellipse are at (1, 5) and (1, –5) and (3, 0) is a point on the ellipse, what is the ellipse equation? (x^2)/(2^2) + (y – 1)^2/(5^2) = 1 (x – 1)^2/(2^2) + (y^2)/(5^2) = 1 (x^2)/(2^2) + (y^2)/(5^2) = 1 (x – 1)^2/(5^2) + (y^2)/(2^2) = 1

OpenStudy (anonymous):

@freckles or @welshfella

OpenStudy (anonymous):

aren't vertices the points

OpenStudy (anonymous):

@freckles

OpenStudy (michele_laino):

If we make this traslation: \[\left\{ \begin{gathered} x = X + 1 \hfill \\ y = Y \hfill \\ \end{gathered} \right.\] then the equation of our ellipse, will be: \[\frac{{{X^2}}}{{{a^2}}} + \frac{{{Y^2}}}{{{b^2}}} = 1\] 

OpenStudy (michele_laino):

what is a?

OpenStudy (anonymous):

2

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

sorry I have made an error, the right equation is: \[\frac{{{X^2}}}{{{b^2}}} + \frac{{{Y^2}}}{{{a^2}}} = 1\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so a is 5^2

OpenStudy (anonymous):

or 5

OpenStudy (michele_laino):

what is b?

OpenStudy (anonymous):

2?

OpenStudy (michele_laino):

starting from the old coordinates, x=3, y=0, we get the new coordinates X, Y using the trslation above, namely: X=2, Y=0 so substituting into the formula of the ellipse, we get: (4/b^2)+(0^2/25)=1 so b=2

OpenStudy (michele_laino):

traslation*

OpenStudy (michele_laino):

then the equation of our ellipse, in the XOY system is: (x^2/4)+(Y^2/25)=1

OpenStudy (michele_laino):

\[\frac{{{X^2}}}{4} + \frac{{{Y^2}}}{{25}} = 1\]

OpenStudy (anonymous):

so its C

OpenStudy (michele_laino):

now, we have to rewrite that equation using the old coordinates, x, and y

OpenStudy (michele_laino):

so we have to replace X with x-1 and Y with y. So, what do you get?

OpenStudy (anonymous):

(x – 1)^2/(2^2) + (y^2)/(5^2) = 1

OpenStudy (michele_laino):

please you have to keep in mind the traslation above. So yes! that's right!

OpenStudy (anonymous):

so its B

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thanks